OpenStudy (anonymous):

Choose the correct simplification of the expression (xyz2)4. (all of the numbers are exponents) x5y5z6 x5y5z8 xyz16 x4y4z8

OpenStudy (zehanz):

\[(xyz^2)^4 =xyz^2xyz^2xyz^2xyz^2=xxxxyyyyz^2z^2z^2z^2 = ...\] It is really a matter of counting...after which you'll see the right answer!